The compression factor (compressibility factor) for one mole of a van der Waal’s at 0oC and 100 atmospheric pressure is found to be 0.5. Assuming that the volume of a gas molecule is negligible, calculate the van der Waals constant a.     

Grade:upto college level

2 Answers

Deepak Patra
9 years ago
We know that, Compressibility factor, Z = PV/RT
0.5 = 100 *V/0.082 *273
∴ V = 0.1117L
NOTE : Further when volume of a gas molecule is negligible, van der Waal’s equation becomes
(P + a/V2) (V - 0) = RT
Or PV = RT – a/V or a = RTV – PV2
Substituting the values
A = (0.082 * 0.1119 *273) – (100 * 0.1119 * 0.1119)
= 1.253 atm L2 mol-2
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
We know that, Compressibility factor, Z = PV/RT
0.5 = 100 *V/0.082 *273
∴ V = 0.1117L
NOTE : Further when volume of a gas molecule is negligible, van der Waal’s equation becomes
(P + a/V2) (V – 0) = RT
Or PV = RT – a/V or a = RTV – PV2
Substituting the values
a = (0.082 * 0.1119 *273) – (100 * 0.1119 * 0.1119)
= 1.253 atm L2 mol-2
 
Thanks and regards,
Kushagra